OpenStudy (idealist10):

Solve for v: v^2+3v+2=Cx

OpenStudy (idealist10):

@phi

OpenStudy (phi):

what is Cx ?

geerky42 (geerky42):

Subtract both sides by Cx and you have \( v^2+3v+2-Cx=0\) Now you can apply quadratic formula, \(\dfrac{-b\pm\sqrt{b^2-4ac}}{2a}\), with \(c=2-Cx\)

OpenStudy (idealist10):

Thank you!